OpenStudy (anonymous):

Solve for the variable in this expression. 3a/7 + 2/3 = 7/9

OpenStudy (texaschic101):

3a/7 + 2/3 = 7/9 --- subtract 2/3 from each side 3a/7 = 7/9 - 2/3 3a/7 = 7/9 - 6/9 3a/7 = 1/9 -- multiply by 63 27a = 7 a = 7/27

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (texaschic101):

anytime :)